HIGH RESOLUTION INFRARED SPECTRA OF CARBON DIOXIDE SOLVATED WITH HELIUM ATOMS

JIAN TANG AND \underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

Infrared spectra of He_{N}-CO_{2} clusters with N up to about 20 have been studied in the region of the CO_{2} ν_{3} fundamental band (2350 cm^{−1}) using a tunable diode laser spectrometer and pulsed supersonic jet source with cooled (> −150 C) pinhole or slit nozzles and high backing pressures (< 40 atm). Compared to previous studies of He_{N}-OCS [1] and -N_{2}O [2] clusters, the higher symmetry of CO_{2} results in simpler spectra but less information content. The binary complex, He-CO_{2}, was studied previously by Weida et al. [3]. With increasing cluster size, N = 2 to 17, we observe discrete rotation-vibration transitions (R(0), P (2), R(2)) whose analysis yields the variation of the band origin and B rotational constant over this size range. The vibrational origin variation is very similar to He_{N}-OCS, with an initial blue shift up to N = 5, followed by a monotonic red shift, consistent with a model where the first 5 He atoms fill a ring around the equator of the molecule, forcing subsequent He atom density to locate closer to the ends. The B value initially drops as expected for a normal molecule, reaching a minimum for N = 5. Its subsequent rise for N = 6 to 11 can be interpreted as the transition from a normal (though floppy) molecule to a quantum solvation regime, where the CO_{2} molecule starts to rotate separately from the He atoms. For N > 13, the B value is approximately constant with a value about 17% larger than that measured in much larger helium nanodroplets [4]. Very recent quantum Monte Carlo simulations by Mezzacapo and Moroni are in excellent agreement with these experimental results [5].

References

[1] J. Tang, Y. Xu, A.R.W. McKellar, and W. Jäger, Science, 297, 2030 (2002).

[2] Y. Xu, W. Jäger, J. Tang, and A.R.W. McKellar, Phys. Rev. Lett., 91, 163401 (2003).

[3] M.J. Weida, J.M. Sperhac, D.J. Nesbitt, and J.M. Hutson, J. Chem. Phys., 101, 8351 (1994).

[4] K. Nauta and R.E. Miller, J. Chem. Phys., 115, 10254 (2001).

[5] J. Tang, A.R.W. McKellar, F. Mezzacapo, and S. Moroni, Phys. Rev. Lett., in press (2004).